RHIC status -- Rotator polarity switch]

From: Flemming Videbaek (videbaek@rcf.rhic.bnl.gov)
Date: Sun Apr 27 2003 - 13:45:10 EDT

  • Next message: Apache: "Shift report 20030428 00:00-08:00"
    Dear Collaborator,
    this weekend has turned out to be less productive for experiment
    due to problems - note below in e-mail from A.Drees. The overall
    plan is still to go into physics mode tomorrow (though the luminosity
    is still dismal in my opnion).
    With the opening now+ retuning I do not expect beam before late.
    
    regards
    
    Flemming
    
    Hey,
    
    (yup, it's me again ;)). The good news first: both beams are polarized
    above 30% at store even with rotators on :). The bad news: it's still
    radial polarization at PHENIX, indicating a wrong polarity of the
    rotators :(. The PS group is switching the polarity and therefore an
    access is granted. PP2PP and STAR are going in under Controlled Access
    as well.  There will be no work with beam before about 16:00 today
    (estimated time for the switch:  approx. 4 h, access started 11:30). Don
    Bruno and George Ganetis are  checking from home.
    
    Unfortunately, after the polarity switch the rotator ramp needs to be
    re-commissioned. This will take a few ramps and work will start later
    this afternoon. I will keep you posted about the progress.
    
    Enjoy this sunny (finally!!) afternoon.
    
    	Angelika (& Adam)
    
    
    -- 
    Flemming Videbaek
    Brookhaven National Laboratory
    Physics Department
    tlf 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    


    This archive was generated by hypermail 2.1.5 : Sun Apr 27 2003 - 13:45:46 EDT